Determine the difference quotient for f(x) = x² + 3x

Answer by Edwin McCravy(20055)



f(x) = x² + 3x

The difference quotient is 

f(x+h)-f(x)
     h

First we find f(x+h) separately:

f(x+h) = (x+h)² + 3(x+h)

f(x+h) = (x+h)(x+h) + 3(x+h)

Multiply the parentheses out and simplify:

f(x+h) = x² + 2hx + h² + 3x + 3h

We put that in place of f(x+h)
and we put (x² + 3x) in place of f(x) in

f(x+h)-f(x)
     h

being sure to put parentheses around (x² + 3x) since
it has a - sign before it:

x² + 2hx + h² + 3x + 3h - (x² + 3x)
                 h 

Next we remove those parentheses we just put around

x² + 2hx + h² + 3x + 3h - x² - 3x
                 h

Notice that all terms cancel except the terms that contain h:

2hx + h² + 3h
      h

Next we factor h out of the top:

h(2x + h + 3)
      h

Finally we cancel the h's

h(2x + h + 3)
      h

2x + h + 3
          
That's it!

[Sometimes "∆x" is used instead of "h". if your course uses "∆x"
then "∆x" must be treated just as "h" above, and never separated
or considered as  "∆" times "x".  Ignore this if your book uses h.]

Edwin